OpenStudy (anonymous):

I have absolutely no clue how to do this problem, will some one please help explain it to me?

OpenStudy (anonymous):

This is what it looks like.

OpenStudy (misty1212):

hi!! no one can do it

OpenStudy (misty1212):

because no one has any idea what T.A is!!

OpenStudy (anonymous):

it means total area

OpenStudy (misty1212):

lolol math teachers can abbreviate anything good morning class gmc

OpenStudy (anonymous):

the diagram makes no sense the 3 is as big as the 6

OpenStudy (anonymous):

if it is supposed to be a square on the bottom, then the area of that square is \(6^2=36\)

OpenStudy (anonymous):

but the 3" throws it off

OpenStudy (anonymous):

not clear is that is supposed to be half of the upper side or all of it the diagram is really not right at all

OpenStudy (anonymous):

See my problem? haha I copied the diagram word for word and none of it makes sense!

OpenStudy (anonymous):

can you post a screen shot of the diagram?

OpenStudy (anonymous):

Some how i am suppose to make an eqation

OpenStudy (anonymous):

yeah i see that, but since the diagram is nonsense you cannot do it right to the 'bad math" pile for this one

OpenStudy (anonymous):

that link may not work since it is on my account

OpenStudy (misty1212):

nah capture it

OpenStudy (anonymous):

That is the picture. The question before it asked me to find the lateral area. I'm pretty sure it is 54 sq. in. but I'm not 100% confident on my answer

OpenStudy (anonymous):

Maybe it's a formula i am putting together? although that isn't clear it makes more sense

OpenStudy (misty1212):

ooh maybe that is supposed to be a 3 dimensional figure?

OpenStudy (anonymous):

Yeah, i have to find the volume afterwards so that makes sense with it being cubed and all

OpenStudy (misty1212):

yikes picture is still weird but it looks like you have two sides of area \(6\times 3=18\) for a total of 36 on the sides

OpenStudy (misty1212):

then we need the area of the two triangles

OpenStudy (misty1212):

which i think they are indicating are equilateral, so the are is \[\frac{\sqrt3 a^2}{4}\]

OpenStudy (misty1212):

you get \[\frac{9\sqrt3}{2}\] the 2 in the denominato rbecause you have two triangles

OpenStudy (anonymous):

Makes a little sense, but it doesn't fit into the question, so it loses me once again

OpenStudy (anonymous):

they have a little square root template there, so why not?

OpenStudy (anonymous):

and don't forget the bottom it is \(6\times 3\) also

OpenStudy (anonymous):

I have to drag the numbers into the right places. If i put it like that equation it doesn't work

OpenStudy (anonymous):

god bless on line systems

OpenStudy (misty1212):

ikr, real homework in a real class, not as bad as it once seemed!!

OpenStudy (anonymous):

It's ridiculous. I failed geometry, so the school makes me do geometry... by myself, no teacher. Yes, that will do me wonders!